1.forty horses are bought for Rs.15000 the average cost of fifteen of them is Rs.300 what is the coaverage st of other?
2.the average of three numbers is 8, that of the first two is 11.find the third number?

Answer by Cromlix(4381)

Total cost of 40 horses = Rs 15000
Average cost of 15 of the horses = Rs 300
Total cost of 15 horses = 300 * 15 = Rs 4500
Total cost of other 25 horses = 15000 - 4500
= Rs 10500.
Average cost for 25 horses = 10500/25 = Rs 420.
.....................................
Average of three numbers is 8
Total of three numbers = 3 * 8 = 24
The average of the two numbers = 11
Total of two numbers = 2 * 11 = 22
Third number = 24 - 22 = 2
Hope this helps.
:-)